OpenStudy (anonymous):

Solve the equation log2x+log(x-1)=log6x for x: A. There are no solutions. B. x = 0,4 C. x = 4 D. x = -1/3

OpenStudy (anonymous):

log[(2x)(x-1)]=log[6x] (2x)(x-1)=6x x-1=3 x=3+1 x=4 note: for log funtions, x must be always greater than 0